OpenStudy (anonymous):

on a scatter plot of data, a trend line passes through the points (13,12) and (11,10). Describe the correlation. -positive correlation -no correlation -negative correlation -none of the above

OpenStudy (anonymous):

slope = 12-10 / 13-11 = 2/2 = 1 slope is positive positive correlation